# Non-Heisenbergian Quantum Mechanics

### v2

@@ III Modified uncertainty relation
where $\xi_{j}\equiv\frac{\partial^{j}}{\partial p^{j}_{i}}f(|\bm{p}-\bm{k}|)|_{\bm{p}=\bm{k}}$, and $\psi(\bm{x})$ is defined using momentum wave function as $\psi(\bm{x})\equiv(2\pi\hbar)^{-3/2}\int\tilde{\psi}(\bm{p})e^{i\bm{x}.\bm{p}/\hbar}d^{3}p$, as usual. Note that, the $\psi(\bm{x})$ should not be interpreted as a ”wave function” in the usual sense, because without a position operator and position eigenstates, $\langle\bm{x}|\psi\rangle$ cannot be defined. Moreover, in the general case, the position probability density is given by eq. (6), not $|\psi(\bm{x})|^{2}$.
The normalization condition, $\int\rho(\bm{x})d^{3}x=1$, ensures that $\xi_{0}=(2\pi\hbar)^{-3}$. Moreover, using the fact that $f$ is a positive kernel, it is easy to prove that $f(x)\leq f(0)$ Wells and Williams (2012), which ensures that $\xi_{1}=0$ and $\xi_{2}\leq 0$. Therefore, if $l_{0}$ is defined as $l_{0}=\sqrt{-\xi_{2}(2\pi\hbar)^{3}\hbar^{2}}$, it is easy to see that,
@@
where $l_{0}$ is a positive constant with length dimension. Therefore, it is clear that,
@@
It should be noted that the assumption that an arbitrary sharp localized state exists, $l_{0}=0$, determines the kernels $f$ uniquely as a constant function $f(|\bm{p}-\bm{k}|)=f(0)=(2\pi\hbar)^{-3}$ Kijowski (1974), which means that the equation (6) reduces to $\rho=|\psi|^{2}$. Moreover, if $f=f(0)$, the equation (7) reduces to the usual Schrödinger probability current, i.e. $\bm{J}=\frac{\hbar}{m}\Im[\psi^{*}\nabla\psi]$. In fact, in this case, the current formalism reduces to the standard Heisenberg quantum theory.
@@
$\int(x_{i}-\langle x_{i}\rangle)^{2}|\psi(\bm{x})|^{2}d^{3}x\int(p_{i}-\langle p_{i}\rangle)^{2}|\tilde{\psi}(\bm{p})|^{2}d^{3}p\geq\hbar^{2}/4,$
@@
This modified uncertainty relation is the same as the common generalized uncertainty relation at the first order, $\Delta x_{i}\Delta p_{i}\geq\hbar/2(1+2l_{0}^{2}(\Delta p_{i})^{2}/\hbar^{2})$, which appears in a wide range of approaches to quantum gravity Amati et al. (1989); Maggiore (1993a); Scardigli (1999); Das and Vagenas (2008); Hossenfelder (2013). Often, such modification is seen as a manifestation of a deformed canonical commutator Maggiore (1993b); Kempf et al. (1995); Fadel and Maggiore (2022) which has been related to a deformation of space-time symmetry Fadel and Maggiore (2022); Maggiore (1994); Bosso et al. (2023a). However, it is important to emphasize that we did not assume any modified symmetry or deformed commutation relation. The general form of quantum probability density (6) and the modified uncertainty relation (9) are derived from the Galilean symmetry.
In the quantum gravity literature, it is often assumed that the value of $l_{0}$ is near the Planck scale ($l_{p}=10^{-35}$ m). Nonetheless, it is worth mentioning that some theories with varying gravitational coupling, such as those with additional dimensions Arkani-Hamed et al. (1998); Antoniadis et al. (1998); Hossenfelder et al. (2003), imply a larger effective Planck length. In our framework, $l_{0}$ is a free parameter, and we do not presuppose any specific value for it. In principle, the value of $l_{0}$ should be determined from a direct comparison of the theoretical result with experimental data Pikovski et al. (2012).
IV Single particle in an external potential
To find an experimental bound on $l_{0}$, we need to extend the model to non-free particles. Firstly, we consider a single particle in the presence of an external potential and then the model is extended for an interacting many-particle system. For a particle in an external potential, we assume that the position probability density is given by a POVM as $\rho_{t}(\bm{x})=\langle\psi_{t}|\hat{Q}_{\bm{x}}|\psi_{t}\rangle$, and the time evolution of the state is given by the following Schrödinger equation:
@@ IV Single particle in an external potential
where $\hat{H}=\hat{\bm{P}}^{2}/2m+\hat{V}$ and the potential operator, associated with the classical scalar potential $V(\bm{x})$, is naturally defined as follows:
@@
Note that, when $\hat{Q}_{\bm{x}}$ are projection operators, this definition reduces to the standard potential operator, i.e., $\hat{V}=\int V(\bm{x})|\bm{x}\rangle\langle\bm{x}|d^{3}x$. In the general case, $l_{0}\neq 0$, this definition is conceptually consistent with our definition of position probability density, eq.(6), due to the fact that this definition ensures
@@
It is easy to show that, the same as in the free particle case, the Galilean symmetry implies that $\langle\bm{p}|\hat{Q}_{\bm{x}}|\bm{k}\rangle=f(|\bm{p}-\bm{k}|)e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}.$ Therefore, in the momentum representation, the modified Schrödinger equation (10) reads
@@
where $\tilde{U}(\bm{p}-\bar{\bm{p}})=f(|\bm{p}-\bar{\bm{p}}|)\int V(\bm{x})e^{i(\bm{p}-\bar{\bm{p}}).\bm{x}/\hbar}.$ The equations (6) and (13) together ensure the local probability conservation—they lead to the continuity equation, $\partial\rho/\partial t+\nabla.\bm{J}=0,$ in which the probability current $\bm{J}$ is given by
$\bm{J}(\bm{x})=\int\left(\alpha(\bm{p}+\bm{k})+\beta(\bm{p}-\bm{k})\right)e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}d^{3}\bm{p}d^{3}\bm{k},$
@@
$\beta=\frac{f(|\bm{p}-\bm{k}|)}{(\bm{p}-\bm{k})^{2}}\times$
$\left(\!\tilde{\psi}^{*}(\bm{k})\!\!\int\!\!\tilde{U}(\bm{p}-\bar{\bm{p}})\tilde{\psi}(\bar{\bm{p}})d^{3}\bar{\bm{p}}\!-\!\tilde{\psi}(\bm{p})\!\!\int\!\!\tilde{U}^{*}\!(\bm{k}-\bar{\bm{k}})\tilde{\psi}^{*}(\bar{\bm{k}})d^{3}\bar{\bm{k}}\right)\!\!.$
It is easy to check that the given probability current is correctly transformed under Galilean transformations Brown and Holland (1999). In principle, this probability current may be experimentally probed using the measurement of arrival time distribution Ali et al. (2003, 2007); Vona et al. (2013); Tumulka (2022); Roncallo et al. (2023); Ayatollah Rafsanjani et al. (2023); Kazemi and Hosseinzadeh (2023).
The only difference between the Eq.(13) and the usual Schrödinger equation in momentum representation is the existence of the factor $f$, which can be effectively considered as deformation of the potential. However it is important to remark that, the physical content of this equation is different from the usual Schrödinger equation with a deformed potential, due to the presence of the factor $f$ in the definition of the position probability density, Eq.(6). Nonetheless, the equations (6) and (13) together lead to
@@
where $\langle\nabla V\rangle$ is naturally defined as $\int\nabla V(\bm{x})\rho(\bm{x})d^{3}x$—which is exactly same as the result in standard quantum theory.
@@
$E\tilde{\psi}(\bm{p})=\frac{\bm{p}^{2}}{2m}\tilde{\psi}(\bm{p})-\frac{m\omega^{2}}{2}\hbar^{2}\nabla^{2}_{\bm{p}}\tilde{\psi}(\bm{p})+\frac{d}{2}ml_{0}^{2}\omega^{2}\tilde{\psi}(\bm{p}),$
which is the same as the standard equation except for the last term. Therefore, in the momentum representation, the harmonic stationary states are the same as standard ones, while the eigen-energies are given by
@@
Here $n=(n_{1},..,n_{d})$ and $n_{i}$ are integer numbers. Note that, although the energy eigenstates are the same as energy eigenstates in usual quantum theory and lead to the same momentum distribution, the position probability density is different from the usual, due to the presence of $f$ kernel in Eq.(6). Moreover, from Eq.(14), it is clear that the time evolution of the position expectation value $\langle x\rangle$ is the same as in standard quantum theory, and is consistent with classical motion in the classical limit—the fact that is in sharp contrast with the results of deformed-algebra formalisms, which are used to experimental probing of the deformed commutation relation Bawaj et al. (2015); Kumar and Plenio (2020). Nonetheless, in our model, the higher-order moments of the position probability density $\langle x^{m}\rangle$ and their time evolution are different from usual cases. To probe these effects some clever experimental design is needed: For example, see the optomechanical setups that are proposed to probe the effect of minimal position uncertainty near the Planck scale Pikovski et al. (2012); Kumar and Plenio (2018); Chevalier et al. (2022). Detailed study of such setups in our non-Heisenbergian quantum theory is beyond the scope of the present work and are left for future studies. Nonetheless, as a first estimation, using Eq.s (15), it is essay to show that
@@
which could be used to estimate an upper limit on $l_{0}$  [^2] Note that, although the potential (and thus the Hamiltonian) is defined up to a constant $C$, the inequality (16) is independent of this constant. If one uses a constantly shifted Hamiltonian, $\hat{H}^{\prime}=\hat{H}+C$, although the Hamiltonian eigenvalues are shifted, the inequality (16) remains correct.. In Refs. Marin et al. (2013, 2014), an experimental limit for the modal energy, i.e., the right-hand side of (16), has been reported by analyzing the residual motion of the first longitudinal mode of the AURIGA gravitational wave bar detector, which is cooled down to sub-millikelvin Vinante et al. (2008). The measured modal energy, $E_{\text{exp}}\!=\!1.3\!\times\!10^{26}$, leads to an upper limit for the minimal position uncertainty as $l_{0}<3.8\times 10^{16}l_{P}$ Marin et al. (2013). Nevertheless, it should be noted that the direct application of the one-particle theory to the center of mass of a macroscopic object with a very large mass, such as the AURIGA detector with $m=10^{13}M_{p}$, may be questionable—at least, the value of $l_{0}$ may not be same for all objects Bosso et al. (2023b); Chevalier et al. (2022). In principle, for a more accurate consideration, a multi-particle extension of our framework is needed. In the next section, we will develop such an extension, and based on it, we obtain a more justifiable upper limit on $l_{0}$ using the observed Hydrogen atom spectrum.
V Interacting particles system
In this section, we extend the theory to multi-particle systems. For simplicity, we only consider a two-particle system; however, extending the results to an $N$-particle system is straightforward. The Hilbert space of the system is considered, as usual, as the tensor product of one-particle Hilbert spaces. The probability density of a two-particle system in configuration space is given by
@@ V Interacting particles system
where, $\{\hat{Q}^{t}_{\bm{x}_{1},\bm{x}_{2}}|(\bm{x}_{1},\bm{x}_{2})\in\mathbb{R}^{6}\}$ is a POVM in the two-particle Hilbert space. Similar to the single-particle theory, the non-relativistic space-time symmetries lead to the following commutation relations:
@@
where, $\bm{x}=(\bm{x}_{1},\bm{x}_{2})$. It is easy to check that the above commutation relations lead to $\langle\bm{k}_{1},\bm{k}_{2}|\hat{Q_{\bm{0}}^{0}}|\bm{p}_{1},\bm{p}_{2}\rangle=f(|\bm{p}_{1}-\bm{k}_{1}|,|\bm{p}_{2}-\bm{k}_{2}|)$, where $f$ is a positive semi-definite kernel in $\mathbb{R}^{6}$. If, as in usual quantum theory, we assume that the separable multi-particle states lead to a separable two-particle probability density—i.e., $|\psi\rangle=\prod_{i}|\psi_{i}\rangle\Rightarrow\rho(\bm{x})=\prod_{i}\rho_{i}(\bm{x}_{i})$—then, we have
@@
where $f_{1}$ and $f_{2}$ are radial-positive-kernels in $\mathbb{R}^{3}$. Therefore, the two-particle probability density in configuration space is given by
@@
where $\bm{x}=(\bm{x}_{1},\bm{x}_{2})$, $\bm{p}=(\bm{p}_{1},\bm{p}_{2})$, and $\bm{k}=(\bm{k}_{1},\bm{k}_{2})$. Note that we consider a system made of two different particles, which, in principle, could have different quantum behaviors characterized by different kernels.
@@
Note that, when $\hat{Q}_{\bm{x}}$ are projection operators, this definition reduces to the standard potential operator—i.e., $\hat{V}=\int V({\bm{x}_{1},\bm{x}_{2}})|{\bm{x}_{1},\bm{x}_{2}}\rangle\langle{\bm{x}_{1},\bm{x}_{2}}|d^{3}x_{1}d^{3}x_{2}$. The Galilean symmetry limits the possible forms of the interaction potential—from $[\hat{\bm{K}},\hat{H}]=i\hbar\hat{\bm{P}}$ and Eq.(20), or from $[\hat{\bm{P}},\hat{H}]=0$ and Eq.(19), it is easy to see that,
@@
Moreover, $[\hat{\bm{J}},\hat{H}]=0$ together with Eq.(21) lead to
@@
The equations (22) and (23) imply that the potential must be a function of $|\bm{x}_{2}-\bm{x}_{1}|$, which is the same condition that arises from Galilean symmetry in conventional quantum mechanics Weinberg (2015).
@@
$E\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})$ $=$ $\left(\frac{\bm{p}^{2}_{1}}{2m_{1}}+\frac{\bm{p}^{2}_{2}}{2m_{2}}\right)\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})$
@@
where, $\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})=\langle\bm{p}_{1},\bm{p}_{2}|\Psi\rangle$, and
$\tilde{U}\!\!=\!\!f_{1}(|\bm{p}_{1}|)f_{2}(|\bm{p}_{2}|)\!\!\!\int\!\!V(|\bm{x}_{1}-\bm{x}_{2}|)e^{\frac{i}{\hbar}(\bm{p}_{1}.\bm{x}_{1}+\bm{p}_{2}.\bm{x}_{2})}d^{3}\bm{x}_{1}d^{3}\bm{x}_{2}.$
It is well-known that, in the case of two interacting particles, working with the total momentum $\bm{p}_{c}=\bm{p}_{1}+\bm{p}_{2}$ and the relative momentum $\bm{p}_{r}=(m_{1}\bm{p}_{2}-m_{2}\bm{p}_{1})/(m_{1}+m_{2})$ is useful. After separation of the external and internal degrees of freedom, $\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})=\tilde{\psi}_{c}(\bm{p}_{c})\tilde{\psi}_{r}(\bm{p}_{r})$, the two-body problem can be split into $E_{c}\psi_{c}(\bm{p}_{c})=\frac{\bm{p}^{2}_{c}}{2M}\psi_{c}(\bm{p}_{c})$, and
@@
where $\tilde{U}_{r}(\bm{p})=(2\pi\hbar)^{3}f_{1}(|\bm{p}|)f_{2}(|\bm{p}|)\int V(|\bm{x}|)e^{i\bm{p}.\bm{r}/\hbar}d^{3}\bm{r}$, $E=E_{c}+E_{r}$, $M=m_{1}+m_{2}$, $\mu=m_{1}m_{2}/(m_{1}+m_{2})$, and $\bm{r}$ represent the relative coordinate, $\bm{x}_{2}-\bm{x}_{1}$. Using the Taylor expansion of $f_{i}$, in the first-order approximation, Eq.(25) reduces to
$E\psi_{r}(\bm{r})\!=\!-\frac{\hbar^{2}}{2\mu}\!\nabla^{2}\psi_{r}(\bm{r})\!-(1+\frac{l_{1}^{2}+l_{2}^{2}}{2}\nabla^{2})V(r)\psi_{r}(\bm{r}),$
@@
in which $a_{0}=4\pi\epsilon_{0}\hbar^{2}/\mu e^{2}$ is the Bohr radius. Comparing this result with the high-precision experimental data of hydrogen atom spectroscopy can lead to an upper bound for the minimal position uncertainties. To achieve this, we focus on the extremely narrow $1S-2S$ two-photon transition. In recent years, the frequency of the $1S-2S$ transition has been accurately measured using a cesium clock with a relative uncertainty of $4.5\times 10^{-15}$ Matveev et al. (2013); Parthey et al. (2011). This relative uncertainty leads to the following upper bound on the minimal position uncertainties:

### v4

@@ 3 Modified uncertainty relation
where $\xi_{j}\equiv\frac{\partial^{j}}{\partial p^{j}_{i}}f(|\bm{p}-\bm{k}|)|_{\bm{p}=\bm{k}}$, and $\psi(\bm{x})$ is defined using momentum wave function as $\psi(\bm{x})\equiv(2\pi\hbar)^{-3/2}\int\tilde{\psi}(\bm{p})e^{i\bm{x}.\bm{p}/\hbar}d^{3}p$, as usual. Note, however, that $\psi(\bm{x})$ should not be interpreted as a "wave function" in the usual sense, because without a position operator and position eigenstates, $\langle\bm{x}|\psi\rangle$ cannot be defined. Moreover, in the general case, the position probability density is given by Eq. (6), not $|\psi(\bm{x})|^{2}$.
The normalization condition, $\int\rho(\bm{x})d^{3}x=1$, ensures that $\xi_{0}=(2\pi\hbar)^{-3}$. Moreover, using the fact that $f$ is a positive kernel, it is easy to prove that $f(x)\leq f(0)$ [49], which ensures that $\xi_{1}=0$ and $\xi_{2}\leq 0$. Therefore, if $l_{0}$ is defined as $l_{0}=\sqrt{-\xi_{2}(2\pi\hbar)^{3}\hbar^{2}}$, it is straightforward to see that
@@
where $l_{0}$ is a positive constant with dimensions of length. Therefore, it is clear that,
@@
It should be noted that the assumption of the existence of arbitrarily sharp localized states, i.e., $l_{0}=0$, uniquely determines the kernel $f$ to be a constant function [52], $f(|\bm{p}-\bm{k}|)=f(0)=(2\pi\hbar)^{-3}$, which implies that equation (6) reduces to $\rho=|\psi|^{2}$. Moreover, if $f=f(0)$, the equation (7) reduces to the usual Schrödinger probability current, i.e. $\bm{J}=\frac{\hbar}{m}\Im[\psi^{*}\nabla\psi]$. In fact, in this case, the current formalism reduces to the standard Heisenberg quantum theory.
@@
$\int\!(x_{i}-\langle x_{i}\rangle)^{2}(p_{i}-\langle p_{i}\rangle)^{2}|\psi(\bm{x})|^{2}|\tilde{\psi}(\bm{p})|^{2}d^{3}xd^{3}p\!\geq\!\frac{\hbar^{2}}{4},\ \$
@@
This modified uncertainty relation is the same as the common generalized uncertainty relation at the first order, $\Delta x_{i}\Delta p_{i}\geq\hbar/2(1+2l_{0}^{2}(\Delta p_{i})^{2}/\hbar^{2})$, which appears in a wide range of approaches to quantum gravity [53, 54, 22, 21, 19]. Often, such modification is seen as a manifestation of a deformed canonical commutator [55, 56, 57] which has been related to a deformation of space-time symmetry [57, 58, 59]. However, it is important to emphasize that we did not assume any modified symmetry or deformed commutation relation. The general form of quantum probability density (6) and the modified uncertainty relation (9) are derived from the Galilean symmetry. A derivation of (an extended version of) the uncertainty relation (9), employing a method similar to ours but with a slightly distinct interpretation and without incorporating constraints from non-relativistic spacetime symmetries, is provided in [60, 61]. See also Appendix B.
In the quantum gravity literature, it is often assumed that the value of $l_{0}$ is near the Planck scale ($l_{p}=10^{-35}$ m). Nonetheless, it is worth mentioning that some theories with varying gravitational coupling, such as those with additional dimensions [62, 63, 64], imply a larger effective Planck length. In our framework, $l_{0}$ is a free parameter, and we do not presuppose any specific value for it. In principle, the value of $l_{0}$ should be determined from a direct comparison of the theoretical result with experimental data [65].
4 Single particle in an external potential
To find an experimental bound on $l_{0}$, we need to extend the model to non-free particles. Firstly, we consider a single particle in the presence of an external potential and then the model is extended for an interacting many-particle system. For a particle in an external potential, we assume that the position probability density is given by a POVM as $\rho_{t}(\bm{x})=\langle\psi_{t}|\hat{Q}_{\bm{x}}|\psi_{t}\rangle$, and the time evolution of the state is governed by the Schrödinger equation:
@@ 4 Single particle in an external potential
where $\hat{H}=\hat{\bm{P}}^{2}/2m+\hat{V}$, and the potential operator, associated with the classical scalar potential $V(\bm{x})$, is naturally defined by
@@
Note that, when $\hat{Q}_{\bm{x}}$ are projection operators, this definition reduces to the standard potential operator, i.e., $\hat{V}=\int V(\bm{x})|\bm{x}\rangle\langle\bm{x}|d^{3}x$. In the general case, where $l_{0}\neq 0$, this definition is conceptually consistent with our definition of the position probability density, Eq. (6), since it ensures that
@@
It is easy to show that, as in the free particle case, Galilean symmetry implies $\langle\bm{p}|\hat{Q}_{\bm{x}}|\bm{k}\rangle=f(|\bm{p}-\bm{k}|)e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}.$ Therefore, in the momentum representation, the modified Schrödinger equation (10) reads
@@
where $\tilde{U}(\bm{p}-\bar{\bm{p}})=f(|\bm{p}-\bar{\bm{p}}|)\int V(\bm{x})e^{i(\bm{p}-\bar{\bm{p}}).\bm{x}/\hbar}d^{3}x.$ Equations (6) and (13) together ensure local probability conservation—they lead to the continuity equation, $\partial\rho/\partial t+\nabla.\bm{J}=0,$ in which the probability current $\bm{J}$ is given by
$\bm{J}(\bm{x})\!=\!\!\int\!\left(\alpha(\bm{p}+\bm{k})\!+\!\beta(\bm{p}-\bm{k})\right)e^{\frac{i}{\hbar}(\bm{p}-\bm{k}).\bm{x}}d^{3}\bm{p}d^{3}\bm{k},$
@@
$\beta=\frac{f(|\bm{p}-\bm{k}|)}{(\bm{p}-\bm{k})^{2}}\left(\!\tilde{\psi}^{*}(\bm{k})\!\!\int\!\!\tilde{U}(\bm{p}-\bar{\bm{p}})\tilde{\psi}(\bar{\bm{p}})d^{3}\bar{\bm{p}}\!-\!\tilde{\psi}(\bm{p})\!\!\int\!\!\tilde{U}^{*}\!(\bm{k}-\bar{\bm{k}})\tilde{\psi}^{*}(\bar{\bm{k}})d^{3}\bar{\bm{k}}\right)\!\!.$
It is easy to check that the given probability current transforms correctly under Galilean transformations [66]. In principle, this probability current may be experimentally probed through measurements of the arrival time distribution [67, 68, 69, 70, 71, 72, 73].
The only difference between the Eq.(13) and the usual Schrödinger equation in the momentum representation is the presence of the factor $f$, which can be effectively considered as a deformation of the potential. However, it is important to remark that the physical content of this equation is different from that of the usual Schrödinger equation with a deformed potential, due to the presence of the factor $f$ in the definition of the position probability density, given in Equation (6). Nonetheless, the equations (6) and (13) together lead to
@@
where $\langle\nabla V\rangle$ is naturally defined as $\int\nabla V(\bm{x})\rho(\bm{x})\,d^{3}x$—which is exactly the same as the result in standard quantum theory.
@@
$E\tilde{\psi}(\bm{p})=\!\left[\frac{\bm{p}^{2}}{2m}-\frac{m\omega^{2}}{2}\hbar^{2}\nabla^{2}_{\bm{p}}+\frac{d}{2}ml_{0}^{2}\omega^{2}\right]\tilde{\psi}(\bm{p}),$
which is the same as the standard equation except for the last term. Therefore, in the momentum representation, the stationary states of the harmonic oscillator are the same as the standard ones, while the eigenenergies are given by
@@
Here, $\bm{n}=(n_{1},\ldots,n_{d})$ and the $n_{i}$ are integers. Note that although the energy eigenstates are the same as those in standard quantum theory and lead to the same momentum distribution, the position probability density differs from the usual case due to the presence of the $f$ kernel in Eq. (6). Moreover, from Eq. (14), it is evident that the time evolution of the position expectation value, $\langle x\rangle$, is the same as in standard quantum theory and is consistent with classical motion in the classical limit—a sharp contrast to the results found in deformed commutation relation frameworks [^3] Using Eq. (14) for a harmonic oscillator, $m\frac{d^{2}\langle x\rangle}{dt^{2}}=-k\langle x\rangle$, it is evident that, in our model—as in standard quantum theory—the position expectation value of a harmonic oscillator follows the classical trajectory precisely. In contrast, the model based on a deformed commutation relation results in anharmonic motion [74, 75], where the oscillation frequency depends on the amplitude. This is a stark contrast that, in principle, provides a possible means of experimentally distinguishing between the two models [74, 76, 77]. Moreover, the deformed commutation relation model leads to a different energy spectrum for a one-dimensional harmonic oscillator [56, 78, 79]: $E_{n}\!=\!\sqrt{1+\!\!\frac{\!m^{2}l_{0}^{4}\omega^{2}}{4\hbar^{2}}}\left(n\!+\!\frac{1}{2}\right)\!\hbar\omega\!+\!\left(\!n^{2}\!+\!n\!+\!\frac{1}{2}\right)\frac{1}{2}ml_{0}^{2}\omega^{2},$ which can be used to distinguish between models.. Nonetheless, in our model, the higher-order moments of the position probability density, $\langle x^{m}\rangle$, and their time evolution differ from those in standard quantum theory. To probe these effects, some clever experimental design is needed. For example, see the optomechanical setups proposed to investigate the effect of minimal position uncertainty near the Planck scale [65, 80, 81]. A detailed study of such setups in our non-Heisenbergian quantum theory is beyond the scope of the present work and is left for future studies. Nonetheless, as a first estimation, using Eq.s (15), it is easy to show that
@@
which could be used to estimate an upper limit on $l_{0}$  [^4] Note that, although the potential (and thus the Hamiltonian) is defined up to an additive constant $C$, the inequality (16) is independent of this constant. If one uses a constantly shifted Hamiltonian, $\hat{H}^{\prime}=\hat{H}+C$, the Hamiltonian eigenvalues are shifted; however, the inequality (16) remains valid.. In Refs. [82, 83], an experimental limit for the modal energy—i.e., the right-hand side of Eq. (16)—has been reported by analyzing the residual motion of the first longitudinal mode of the AURIGA gravitational wave bar detector, which was cooled down to sub-millikelvin temperatures [84]. The measured modal energy, $E_{\text{exp}}\!=\!1.3\!\times\!10^{-26}\text{J}$, leads to an upper bound on the minimal position uncertainty as $l_{0}<3.8\times 10^{16}l_{P}$ [82].
Nevertheless, it should be noted that the direct application of the one-particle theory to the center of mass of a macroscopic object with a very large mass, such as the AURIGA detector with $m=10^{13}M_{p}$, may be questionable—at least, the value of $l_{0}$ may not be the same for all objects [85, 81]. In principle, for a more accurate consideration, a multi-particle extension of our framework is needed. In the next section, we develop such an extension, and based on it, we obtain a more justifiable upper limit on $l_{0}$ using the observed hydrogen atom spectrum.
5 Interacting particles system
In this section, we extend the theory to multi-particle systems. For simplicity, we only consider a two-particle system; however, extending the results to an $N$-particle system is straightforward. As usual, the Hilbert space of the system is taken to be the tensor product of the individual one-particle Hilbert spaces. The probability density of a two-particle system in configuration space is given by
@@ 5 Interacting particles system
where $\{\hat{Q}^{t}_{\bm{x}_{1},\bm{x}_{2}}|(\bm{x}_{1},\bm{x}_{2})\in\mathbb{R}^{6}\}$ is a POVM in the two-particle Hilbert space. Similar to the single-particle theory, the non-relativistic space-time symmetries lead to the following commutation relations:
@@
where $\bm{x}=(\bm{x}_{1},\bm{x}_{2})$. It is easy to check that the above commutation relations lead to $\langle\bm{k}_{1},\bm{k}_{2}|\hat{Q_{\bm{0}}^{0}}|\bm{p}_{1},\bm{p}_{2}\rangle=f(|\bm{p}_{1}-\bm{k}_{1}|,|\bm{p}_{2}-\bm{k}_{2}|)$, where $f$ is a positive semi-definite kernel in $\mathbb{R}^{6}$. If, as in usual quantum theory, we assume that the separable multi-particle states lead to a separable two-particle probability density—i.e., $|\psi\rangle=\prod_{i}|\psi_{i}\rangle\Rightarrow\rho(\bm{x})=\prod_{i}\rho_{i}(\bm{x}_{i})$—then we have
@@
where $f_{1}$ and $f_{2}$ are radial positive kernels in $\mathbb{R}^{3}$. Therefore, the two-particle probability density in configuration space is given by
@@
where $\bm{x}=(\bm{x}_{1},\bm{x}_{2})$, $\bm{p}=(\bm{p}_{1},\bm{p}_{2})$, and $\bm{k}=(\bm{k}_{1},\bm{k}_{2})$. Note that we consider a system consisting of two different particles, which, in principle, could have different quantum behaviors characterized by different kernels.
@@
It should be noted that, when $\hat{Q}_{\bm{x}}$ are projection operators, this definition reduces to the standard potential operator—i.e., $\hat{V}=\int V({\bm{x}_{1},\bm{x}_{2}})|{\bm{x}_{1},\bm{x}_{2}}\rangle\langle{\bm{x}_{1},\bm{x}_{2}}|d^{3}x_{1}d^{3}x_{2}$. The Galilean symmetry limits the possible forms of the interaction potential—from $\hat{\bm{K}},\hat{H}]=i\hbar\hat{\bm{P}}$ and Eq.(20), or from $\hat{\bm{P}},\hat{H}]=0$ and Eq.(19), it is easy to see that,
@@
Moreover, $\hat{\bm{J}},\hat{H}]=0$ together with Eq.(21) lead to
@@
The equations (22) and (23) imply that the potential must be a function of $|\bm{x}_{2}-\bm{x}_{1}|$, which is the same condition that arises from Galilean symmetry in conventional quantum mechanics [46].
@@
$E\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})=\left(\frac{\bm{p}^{2}_{1}}{2m_{1}}+\frac{\bm{p}^{2}_{2}}{2m_{2}}\right)\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})$ (24)
@@
where $\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})=\langle\bm{p}_{1},\bm{p}_{2}|\Psi\rangle$, and
$\tilde{U}(\bm{p}_{1},\bm{p}_{2})=f_{1}(|\bm{p}_{1}|)f_{2}(|\bm{p}_{2}|)$
$\times\!\int V(|\bm{x}_{1}-\bm{x}_{2}|)e^{\frac{i}{\hbar}(\bm{p}_{1}.\bm{x}_{1}+\bm{p}_{2}.\bm{x}_{2})}d^{3}\bm{x}_{1}d^{3}\bm{x}_{2}.$
It is well-known that, in the case of two interacting particles, working with the total momentum $\bm{p}_{c}=\bm{p}_{1}+\bm{p}_{2}$ and the relative momentum $\bm{p}_{r}=(m_{1}\bm{p}_{2}-m_{2}\bm{p}_{1})/(m_{1}+m_{2})$ is useful. After separating the external and internal degrees of freedom, $\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})=\tilde{\psi}_{c}(\bm{p}_{c})\tilde{\psi}_{r}(\bm{p}_{r})$, the two-body problem can be split into $E_{c}\psi_{c}(\bm{p}_{c})=\frac{\bm{p}^{2}_{c}}{2M}\psi_{c}(\bm{p}_{c})$, and
@@
where $\bm{r}$ represents the relative coordinate, $\bm{x}_{2}-\bm{x}_{1}$, $\mu=m_{1}m_{2}/(m_{1}+m_{2})$, $E=E_{c}+E_{r}$, $M=m_{1}+m_{2}$, and
$\tilde{U}_{r}(\bm{p})=(2\pi\hbar)^{3}f_{1}(|\bm{p}|)f_{2}(|\bm{p}|)\int V(|\bm{x}|)e^{i\bm{p}.\bm{r}/\hbar}d^{3}\bm{r}.$
Using the Taylor expansion of $f_{i}$, in the first-order approximation, Eq.(25) reduces to
$\left[E+\frac{\hbar^{2}}{2\mu}\!\nabla^{2}+(1\!+\!\frac{l_{1}^{2}+l_{2}^{2}}{2}\nabla^{2})V(r)\right]\psi_{r}(\bm{r})=0,$
@@
in which $a_{0}=4\pi\epsilon_{0}\hbar^{2}/\mu e^{2}$ is the Bohr radius. Comparing this result with the high-precision experimental data of hydrogen atom spectroscopy can lead to an upper bound for the minimal position uncertainties. To achieve this, we focus on the extremely narrow 1S2S1S-2S two-photon transition. In recent years, the frequency of the 1S2S1S-2S transition has been measured with high accuracy using a cesium clock with a relative uncertainty of $4.5\times 10^{-15}$ [86, 87]. This relative uncertainty leads to the following upper bound on the minimal position uncertainties: